# Learning Diffeomorphism for Image Registration with Time-Continuous Networks using Semigroup Regularization

### v2

@@ 6 Conclusion
In this paper, we proposed SGDIR, a novel learning-based model for diffeomorphic image registration based on time-embedded UNet architecture and using semigroup-based regularization. The time-embedded network allows us to compute deformation and consequently warp images at any time point. The semigroup regularization enables the model to generate diffeomorphic deformations and its inverse. The resulting model is capable of generating smooth warpings from the moving image to the fixed image and vice versa without requiring any integration scheme or using further regularization terms on the smoothness of the deformation. We evaluated our model on four benchmark datasets on 3D medical image registration and demonstrated superiority in various metrics compared to the state-of-the-art optimization and learning-based methods. In a broader perspective, the proposed framework provides a systematic method for addressing problems which can be modeled as an autonomous ODE.

### v3

@@ 6 Conclusion
In this paper, we proposed SGDIR, a novel learning-based model for diffeomorphic image registration with a time-embedded UNet architecture and using a semigroup-based regularization. The time-embedded network allows us to compute deformation and consequently warp images at any time point, and the semigroup regularization enables the model to generate diffeomorphic deformations and its inverse. The resulting model is capable of generating smooth warpings from the moving image to the fixed image and vice versa without requiring any integration scheme or using further regularization terms on the smoothness and inverse consistency. We evaluated our model on four benchmark datasets on 3D medical image registration and demonstrated its superiority in various metrics compared to the state-of-the-art methods. In a broader perspective, the proposed framework provides a systematic method for addressing problems which can be modeled as an autonomous ODE.